OpenStudy (anonymous):

WILL FAN AND MEDAL!!! In a sample of 44 city residents, 11 reported that they take vitamin C supplements daily. If there are 24,888 residents in the city, estimate the number of city residents who take vitamin C supplements daily

OpenStudy (anonymous):

@mathstudent55

OpenStudy (zpupster):

this is a ratio 11/44 = x/24888 11(24888) /44 = x

OpenStudy (hockeychick23):

its 25% of 24,888, so .25(24,888)

OpenStudy (hockeychick23):

the number of residents who take vitamin C supplements daily is 6,222

OpenStudy (zpupster):

yes if you recognize that 11/44 = 1/4 and that is 25% or .25